If 100,000 people are shot every year, out of roughly 300 million Americans - what are the odds of being shot in a lifetime?
Without any real number crunching (because I don't know how), I got one-in-13. I figured that it is one-in-3000 in a given year. Then I asked google, and I found that a person has a one-in-700,000 chance of being struck by lightning in a given year which means a one-in-3000 chance in a lifetime (their math based upon an 80-year lifetime). I divided 700 by 3 and got 233.33, so I divided 3000 by 233.33 in kind and got 12.85 (I rounded down because 80 years is generous). I would guess that being shot would shorten one's life expectancy, so this is far a perfect model...
Am I anywhere remotely close?

Answer by stanbon(75887)

If 100,000 people are shot every year, out of roughly 300 million Americans - what are the odds of being shot in a lifetime?
----
P(being shot) = 100,000/300,000,000 = 100/300,000
P(not being shot) = 2,900,000/300,000,000 = 2900/300,000
======
Odds for being shot = 100:2900 = 1:29
-------------------------------------------
Cheers,
Stan H.
====================